OpenStudy (anonymous):

Find and simplify the difference quotient f(x + h)− f(x)/h for the following function. f(x) = 3/x

OpenStudy (campbell_st):

this is 1st principals... f(x)= 3/x f(x +h) = 3/(x + h) you need a common factor f(x +h) - f(x) = 3x - 3(x + h) ------------ (x +h)(x) hope this makes sense as a start